A father left 1/2 of his estate to his son, 1/3 of his estate to his granddaughter, and the remaining $6,000 to charity. What was his total estate?
Answer by edjones(8007)

1/2 + 1/3
=3/6 + 2/6
=5/6
1- 5/6 = 1/6
.
Let x=the total estate
x/6=6000
x=$36,000
.
Ed